Dr. Dixit joggged a total of 6000 yards in three nights. If each night he increased his distance 440 yard, how far did he jog on the first night?
Answer by ikleyn(52785)

.

Let x be the distance he jogged on the SECOND night.


Then on the first nigh he jogged x-440 yards,
while on the third night he jogged x+440 yards.


In total,  (x-440) + x + (x+440) = 6000,  which gives

    3x = 6000,  x = 6000/3 = 2000.


On the first night, he jogged 2000-440 = 1560 yards.    ANSWER

Solved.